I want to estimate the proportion of adults who like sci-fi movies. How many people should I poll to be 96% confident that I am with 1% of the population proportion?

n=
Answer #1

solution: M.E=1% =0.01 d=1-0.96=0.04 C-961 -0.96, 24/2=2.06 zx/2 2 n- PC-p) ME 2.06 xo.5X0.5 0.01 in= 10609